two airplanes begin 1000 miles apart and fly along lines that intersect at right angles one plane flies an average of 100 miles per hour faster than the other. if the planes meet after 2 hours how fast do the airplanes fly?
Answer by mananth(16949)

two airplanes begin 1000 miles apart and fly along lines that intersect at right angles one plane flies an average of 100 miles per hour faster than the other. if the planes meet after 2 hours how fast do the airplanes fly?
speed of one plane be x mph
other plane speed = x+100 mph
...
distance traveled in 2 hours = 2x & 2*(x+100)miles
(2x)^2+(2*(x+100))^2=1000^2
4x^2+4*(x^2+200x+10000)=1000000
/4
x^2+x^2+200x+10000=250000
2x^2+200x-240000=0
/2
x^2+100x-120000=0
x^2+400x-300x-120000=0
(x+400)(x-300)=0
..
the speeds are 300, 400 mph
m.ananth@hotmail.ca